# The Existence of MacWilliams-Type Identities for the Lee, Homogeneous and Subfield Metric

### v2

@@ 2.2. Group Characters
Let $\mathcal{P}=P_{1}\mid P_{2}\mid\dots\mid P_{m}$ be a partition of a finite abelian group $G$. Let $\widehat{\mathcal{P}}=Q_{1}\mid Q_{2}\mid\dots\mid Q_{\ell}$ denote the partition of $\widehat{G}$ where $\chi,\chi^{\prime}\in\widehat{G}$ lie in the same block $Q_{j}$, for $j=1,\ldots,\ell$, if and only if, for every $i=1,\ldots,m$, it holds
@@
In [19], Gluesing-Luerssen referred to a partition $\mathcal{P}$ as Fourier-reflexive if $\widehat{\widehat{\mathcal{P}}}=\mathcal{P}$, and showed that a partition $\mathcal{P}$ is Fourier-reflexive if and only if $\mathcal{P}$ and $\widehat{\mathcal{P}}$ consist of the same number of blocks. This definition aligns with the Fourier-invariant partitions introduced in [47].
@@
Let $\mathcal{P}=P_{1}\mid P_{2}\mid\dots\mid P_{m}$ be a self-dual partition of a finite abelian group $G$. For $a\in P_{j}$, let $\chi_{a}$ be a character of $G$. If the sum $\sum_{x\in P_{i}}\chi_{a}(x)$ does not depend on the choice of $a\in P_{j}$, we define the Krawtchouk coefficient of $i$ with respect to $j$ as
@@
The existence of the Krawtchouk coefficients depends on the partition of the group, and [47, Lemma 1] shows that the Krawtchouk coefficients are well-defined if and only if the partitions are Fourier-invariant. Additionally, Gluesing-Luerssen in [19] showed that reflexive partitions lead to the MacWilliams identities.
In coding theory, partitions are usually induced by a weight function on the ambient space. Given a finite chain ring $\mathcal{R}$ and an additive weight $\operatorname{wt}$ on $\mathcal{R}^{n}$, we can partition $\mathcal{R}^{n}$ into tuples of the same weight. That is, we consider a partition $\mathcal{P}=P_{0}\;|\;\ldots\;|\;P_{\left|\,{\operatorname{Im}(\operatorname{wt})}\,\right|}$ where
(35) $P_{i}\coloneqq\{x\in\mathcal{R}^{n}\;|\;\operatorname{wt}(x)=i\}.$
Example 18.
Let $\mathcal{P}=P_{0}|\ldots|P_{n}$ be the Hamming weight partition of $\mathbb{F}_{q}^{n}$. That is, for any $0\leq i\leq n$,
(36) $P_{i}=\{x\in\mathbb{F}_{q}^{n}\;|\;\operatorname{wt}_{\mathsf{H}}(x)=i\}\ .$
Given $a\in P_{j}$, let $\chi_{a}$ be a character over $\mathbb{F}_{q}^{n}$. The Krawtchouk coefficients of the Hamming partition are well-defined and are described by the following explicit formula [16, Section 4.1],
(37) $K_{j}(i)=\sum_{x\in P_{i}}\chi_{a}(x)=\sum_{k=0}^{j}(-1)^{k}(q-1)^{j-k}\binom{i}{k}\binom{n-i}{j-k}\ .$
Therefore, the Hamming partition of $\mathbb{F}_{q}^{n}$ is Fourier-invariant.
2.3. The MacWilliams Identities over Finite Chain Rings
We finally recap the MacWilliams identities for the Hamming metric and the non-existence results in the Lee, homogeneous and subfield metric in this section. Again, let $\mathcal{R}$ be a finite chain ring and let $\mathrm{wt}\colon\mathcal{R}\longrightarrow\mathbb{Q}$ be an additive weight over $\mathcal{R}$. Furthermore, we consider a linear code $\mathcal{C}\subseteq\mathcal{R}^{n}$. From now on, we will denote by $M_{\mathrm{wt}}$ the maximum value of $\mathrm{wt}$ on $\mathcal{R}$.
The $\mathrm{wt}$-weight distribution of $\mathcal{C}$ is a tuple which specifies the number of codewords of each possible weight $i\in\left\{{0,\ldots,nM_{\operatorname{wt}}}\right\}$, i.e.,
@@ 2.3. The MacWilliams Identities over Finite Chain Rings
where $A_{i}^{\mathrm{wt}}(\mathcal{C}):=\left|\,{\left\{{c\in\mathcal{C}\;|\;\operatorname{wt}(c)=i}\right\}}\,\right|$. If $\mathrm{wt}$ and $\mathcal{C}$ are clear from the context, we will simply write $A_{i}$ instead of $A_{i}^{\mathrm{wt}}(\mathcal{C})$. Moreover, the weight distribution determines the $\mathrm{wt}$-weight enumerator of $\mathcal{C}$:
@@
Theorem 19 (MacWilliams Identities [31, Lemma 2.2]).
@@
where the Krawtchouk coefficients $K_{j}(i)$ are the positive integers given in (37).
Some works on MacWilliams identities investigate partitions of the ambient space. Indeed, it is well-known that if the partition of the ambient space is Fourier-reflexive, the MacWilliams-type identities for this partition exist [19].
@@
In both the cases of Lee and homogeneous metric, as well as for the subfield metric (as shown in [22]), the Krawtchouk coefficients, introduced in Definition 17, are not well-defined as $\sum_{x\in P_{i}}\chi_{a}(x)$ depends on the choice of $a\in P_{j}$. Therefore, we cannot use Theorem 20 to obtain the MacWilliams identities. This leads us to considering a different partition of the ambient space.
In the following, we partition the ambient space based on a natural decomposition stemming from the Lee metric that allows us to partition any ambient space. The considered partition is finer than the weight partition and enables to recover any weight enumerator defined over the ambient space.
We want to emphasize, that this technique was already known to MacWilliams: she used a similar approach to show MacWilliams-type identities for codes over finite fields endowed with the Lee metric.
3. MacWilliams-Type Identities
As shown in Theorems 21 and 22, the MacWilliams identities do not exist for the Lee weight, for the homogeneous or the subfield weight, apart from some trivial cases. As highlighted, the problem mainly lies in the fact that, for the respective weight partitions,
(42) $\sum_{x\in P_{j}}\xi^{\operatorname{Tr}(\left\langle{a,x}\right\rangle)}$
depends on the choice of $a\in P_{i}$, leading to Krawtchouk coefficients which are not well-defined.
In this section, we present MacWilliams-type identities for additive weights. More precisely, we start by showing the identities with respect to a partition suitable for the Lee metric which we refer to as the Lee partition. This approach and the MacWilliams identities resulting from it have already been found by MacWilliams herself in [32] over a finite field where she partitioned the code into codewords of the same Lee decomposition. In her work, MacWilliams represents the finite field $\mathbb{F}_{p^{r}}$ as $\mathbb{F}_{p^{r}}=\left\{{\omega_{0}=0,\pm\omega_{1},\ldots,\pm\omega_{\left\lfloor{p^{r}/2}\right\rfloor}}\right\}$ and refers to the Lee decomposition of a vector $v\in\mathbb{F}_{p^{r}}$ as the tuple whose $i$-th coordinate indicates the number of entries of $v$ equal to either $\omega_{i}\in\mathbb{F}_{p^{r}}$ or $-\omega_{i}\in\mathbb{F}_{p^{r}}$. It should be pointed out that the Lee metric is not properly defined on a non-prime finite field $\mathbb{F}_{p^{r}}$. Hence, over a finite field, this decomposition does not immediately recover the Lee weight of $v\in\mathbb{F}_{p^{r}}$. In [6, 41] similar MacWilliams-type identities for the Lee decomposition have been derived using association schemes. Additionally, it was shown in [19] that partitioning the ambient space into all singletons yields a Fourier-reflexive partition for which the MacWilliams identities exist. In the same paper it was shown that partitions based on the orbits of a subgroup of the automorphism group of the ambient space are also Fourier-reflexive. In [34] the case of symmetrized weight enumerators is discussed. We show that the identities hold for any finite chain ring $\mathcal{R}$ and any additive weight over $\mathcal{R}$. In particular, the identities hold also for the homogeneous and subfield metric.
3.1. Lee partition for Additive Weights
In the following, we consider $\mathcal{R}$ to be any finite chain ring with residue field size $q=p^{r}$, nilpotency index $s$ and $\mathcal{C}\subseteq\mathcal{R}^{n}$ a linear code. Moreover, let $\operatorname{wt}$ be an additive weight over $\mathcal{R}$. In this section, we label the elements of $\mathcal{R}$ as follows:
(43) =$\mathcal{R}=\{\alpha_{0}=0,\ \alpha_{1},\ \dots,\alpha_{M},\ \alpha_{M+1},\dots\ \alpha_{q^{s}-1}\},$
@@ 3.1. Lee partition for Additive Weights
where $M\coloneqq\left\lfloor{\frac{q^{s}}{2}}\right\rfloor$ and $\alpha_{q^{s}-i}$ represents the additive inverse of $\alpha_{i}$ with respect to addition, i.e., $\alpha_{q^{s}-i}=-\alpha_{i}$. We refer to $\alpha_{q^{s}-i}$ as the opposite of $\alpha_{i}$.
Let us introduce the Lee partition of $\mathcal{R}$ where the blocks are formed by an element and its opposite.
@@
Given a finite chain ring $\mathcal{R}$, for any $0\leq i\leq M$ we set $I^{\mathsf{L}}_{i}\coloneqq\{\alpha_{i},-\alpha_{i}\}$ to be the subset of $\mathcal{R}$ containing $\alpha_{i}$ and its opposite. We then define the Lee partition of $\mathcal{R}$ as
(44) $\mathcal{I}^{\mathsf{L}}=I^{\mathsf{L}}_{0}\;|\;I^{\mathsf{L}}_{1}\;|\;\ldots\;|\;I^{\mathsf{L}}_{M}\ .$
The Lee partition $\mathcal{I}^{\mathsf{L}}$ of $\mathcal{R}$ gives rise to a Lee decomposition of a tuple $x\in\mathcal{R}^{n}$ capturing the number of entries equal to ±$\pm\alpha_{i}$ , for all $0\leq i\leq M$.
Definition 24.
Given a finite chain ring $\mathcal{R}$, let $\mathcal{I}^{\mathsf{L}}=I^{\mathsf{L}}_{0}\;|\;\ldots\;|\;I^{\mathsf{L}}_{M}$ be the Lee partition of $\mathcal{R}$. We define the Lee decomposition of an $n$-tuple $x\in\mathcal{R}^{n}$ as the $(M+1)$-tuple $\pi^{\mathsf{L}}(x)=(\pi^{\mathsf{L}}_{0}(x),\ldots,\pi^{\mathsf{L}}_{M}(x))$, where for each $i\in\left\{{0,\ldots,M}\right\}$
(45) $\pi^{\mathsf{L}}_{i}(x)=\left|\,{\left\{{k=1,\ldots,n\;|\;x_{k}\in I^{\mathsf{L}}_{i}}\right\}}\,\right|.$
Therefore, the Lee partition of $\mathcal{R}$ induces a partition of the ambient space $\mathcal{R}^{n}$.
Definition 25.
Let $\mathcal{R}$ be a finite chain ring and $n$ a positive integer. The set of all Lee decompositions in $\mathcal{R}^{n}$ is denoted by
(46) $\mathbb{D}^{\mathsf{L}}_{p,r,s,n}:=\left\{{\pi\in\left\{{0,\ldots,n}\right\}^{M}\,\Big{|}\,\sum_{i=0}^{M}\pi_{i}=n}\right\}.$
Moreover, given $\mathbb{D}^{\mathsf{L}}_{p,r,s,n}=\left\{{\pi^{(0)},\dots,\pi^{(D-1)}}\right\}$, for some positive integer $D$, we define the Lee partition of $\mathcal{R}^{n}$ as
(47) $\mathcal{P}^{\mathsf{L}}=P^{\mathsf{L}}_{\pi^{(0)}}\;|\;P^{\mathsf{L}}_{\pi^{(1)}}\;|\;\ldots\;|\;P^{\mathsf{L}}_{\pi^{(D-1)}}.$
That is, for every $i\in\left\{{0,\ldots,D-1}\right\}$
(48) $P^{\mathsf{L}}_{\pi^{(i)}}\coloneqq\left\{{x\in\mathcal{R}^{n}\;|\;\pi^{\mathsf{L}}(x)=\pi^{(i)}}\right\}.$
Given two Lee decompositions $\pi,\rho\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}$, with the aim of computing the component-wise product of two tuples with Lee decompositions $\pi$ and $\rho$, respectively, we denote the set of all compositions of $\rho$ with respect to $\pi$ by
(49) $\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho):=\Big{\{}t=(t_{0},\ldots,t_{M})\,\Big{|}\,$ $t_{i}=(t_{i0},\ldots,t_{iM})\in\left\{{0,\ldots,\pi_{i}}\right\}^{M}\text{ with }$
(50) $\sum_{j=0}^{M}t_{ij}=\pi_{i}\text{ and }\sum_{i=0}^{M}t_{ij}=\rho_{j},\;\forall i,j\in\left\{{0,\ldots,M}\right\}\Big{\}}.$
Figure 1 illustrates the set $\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)$, i.e., the compositions of $\rho\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}$ with respect to $\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}$.

### v3

@@ 2.3. Partitions and Induced Partitions
The partition of a group $G$ can be extended to a partition of the $n$-fold cartesian product of the group $G$ as follows [8, 18].
Definition 15.
Let $G$ be a finite abelian, nontrivial group and consider a partition $\mathcal{P}=P_{1}\;|\;\cdots\;|\;P_{m}$ of $G$. For $g,g^{\prime}\in G^{n}$, define the equivalence relation $g\sim g^{\prime}$ if $|\left\{{i\in\{1,\ldots,n\}\;|\;g_{i}\in P_{j}}\right\}|=|\left\{{i\in\{1,\ldots,n\}\;|\;g^{\prime}_{i}\in P_{j}}\right\}|$ for all $j\in\{1,\ldots,m\}$. The resulting partition $\mathcal{P}_{\mathrm{sym},n}$ on $G^{n}$ is called the induced symmetrized partition.
@@
Let $\mathcal{P}=P_{1}\mid P_{2}\mid\dots\mid P_{m}$ be a partition of a finite abelian group $G$. Let $\widehat{\mathcal{P}}=Q_{1}\mid Q_{2}\mid\dots\mid Q_{\ell}$ denote the partition of $\widehat{G}$ where $\chi,\chi^{\prime}\in\widehat{G}$ lie in the same block $Q_{j}$, for $j\in\{1,\ldots,\ell\}$, if and only if, for every $i=\in\{1,\ldots,m\}$, it holds
@@
In [18], Gluesing-Luerssen referred to a partition $\mathcal{P}$ as Fourier-reflexive if $\widehat{\!\widehat{\mathcal{P}}}=\mathcal{P}$, and showed that a partition $\mathcal{P}$ is Fourier-reflexive if and only if $\mathcal{P}$ and $\widehat{\mathcal{P}}$ consist of the same number of blocks. This definition aligns with the Fourier-invariant partitions introduced in [25]. Additionally, [18, Theorem 3.3] shows that an induced symmetrized partition $\mathcal{P}_{\mathrm{sym},n}$ is Fourier-reflexive if $\mathcal{P}$ is.
@@
Let $\mathcal{P}=P_{1}\mid P_{2}\mid\dots\mid P_{m}$ and $\mathcal{Q}=Q_{1}\mid Q_{2}\mid\dots\mid Q_{m}$ be two partitions of the finite abelian group $G$. For $a\in Q_{j}$, let $\chi_{a}$ be a character of $G$. If the sum $\sum_{x\in P_{i}}\chi_{a}(x)$ does not depend on the choice of $a\in Q_{j}$, we define the Krawtchouk coefficient of $i$ with respect to $j$ as
@@
Identifying the group $G$ with its character group $\widehat{G}$, the partition $\mathcal{Q}$ induces a partition on $\widehat{G}$. Zinoviev and Ericson [25, Lemma 1] showed that the Krawtchouk coefficients are well-defined if and only if the partitions are Fourier-invariant. Furthermore, the existence of the Krawtchouk coefficients depends on the partition of the group rather than the group itself. In coding theory, partitions are often induced by a weight function on the ambient space. Given a finite chain ring $\mathcal{R}$ and an additive weight $\operatorname{wt}$ on $\mathcal{R}^{n}$, we can partition $\mathcal{R}^{n}$ into tuples of the same weight. That is, we consider a partition $\mathcal{P}^{\operatorname{wt}}=P^{\operatorname{wt}}_{0}\;|\;\cdots\;|\;P^{\operatorname{wt}}_{\left|\,{\operatorname{Im}(\operatorname{wt})}\,\right|-1}$ where
(26) $P^{\operatorname{wt}}_{i}\coloneqq\{x\in\mathcal{R}^{n}\;|\;\operatorname{wt}(x)=i\}.$
2.4. MacWilliams Identities over Finite Chain Rings
We finally recap the MacWilliams identities for the Hamming metric and the non-existence results in the Lee, homogeneous in this section.
Again, let $\mathcal{R}$ be a finite chain ring and let $\mathrm{wt}\colon\mathcal{R}^{n}\longrightarrow\mathbb{Q}$ be an additive weight over $\mathcal{R}$. Furthermore, we consider a linear code $\mathcal{C}\subseteq\mathcal{R}^{n}$. From now on, we will denote by $M_{\mathrm{wt}}$ the maximum value of $\mathrm{wt}$ on $\mathcal{R}$. The $\mathrm{wt}$-weight distribution of $\mathcal{C}$ is a tuple which specifies the number of codewords of each possible weight $i\in\left\{{0,\ldots,nM_{\operatorname{wt}}}\right\}$, i.e.,
@@ 2.4. MacWilliams Identities over Finite Chain Rings
where $A_{i}^{\mathrm{wt}}(\mathcal{C}):=\left|\,{\left\{{c\in\mathcal{C}\;|\;\operatorname{wt}(c)=i}\right\}}\,\right|$. If $\mathrm{wt}$ and $\mathcal{C}$ are clear from the context, we will simply write $A_{i}$ instead of $A_{i}^{\mathrm{wt}}(\mathcal{C})$. Moreover, the weight distribution determines the $\mathrm{wt}$-weight enumerator of $\mathcal{C}$ defined as
@@
Theorem 18 (MacWilliams Identities [1, Lemma 2.2]).
@@
where the $Krawtchouk$ coefficients are given by $K_{j}(i)=\sum_{k=0}^{j}(-1)^{k}(q-1)^{j-k}\binom{i}{k}\binom{n-i}{j-k}$.
@@
Both, the Lee and homogeneous weight enumerator, can be interpreted as symmetrized weight partitions on $\mathcal{R}^{n}$ induced by the weight partition on $\mathcal{R}$ and are not Fourier-reflexive. This motivates the need to consider a different partition of the ambient space.
3. Partition-based MacWilliams Identities
As discussed in the previous section, the MacWilliams identities generally fail for some weight enumerators like, for instance, the Lee weight enumerator and the homogeneous weight enumerator (see Theorems 20 and 21, respectively). However, using partitions different from the weight partition yields MacWilliams-type identities provided that the partition is Fourier-reflexive, as established in [18]. In this section we make use of these results and present for different additive metrics over $\mathcal{R}$ each a symmetrized partition that is Fourier-reflexive (and hence satisfies the MacWilliams-type identities) and allows to retrieve the weight enumerator in the metric considered. Additionally, we show that the partitions presented for each metric are the coarsest such partition. Hence, we refer to such a partition as the coarsest Fourier-reflexive symmetrized partition with respect to the metric considered.
We start by introducing the set-up that we use throughout this section for all metrics. Let $\mathcal{R}$ denote a finite chain ring and $\operatorname{wt}$ a weight function over $\mathcal{R}$ extending additively to $\mathcal{R}^{n}$. As we are looking for a Fourier-reflexive partition that enables recovering the weight enumerator of a code $\mathcal{C}\subseteq\mathcal{R}^{n}$ and since the weight $\operatorname{wt}$ considered is additive, a natural choice is to focus on induced symmetrized partitions of $\mathcal{R}^{n}$ (see Definition 15).
Given a partition $\mathcal{P}=I_{0}\;|\;\cdots\;|\;I_{m}$ on $\mathcal{R}$, we denote by $\mathcal{P}_{\mathrm{sym},n}$ the induced symmetrized partition of $\mathcal{R}^{n}$. The blocks of the symmetrized partitions are defined by the distinct decompositions $\pi$ that an $n$-tuple $x\in\mathcal{R}^{n}$ admits. That is, we define the $\mathcal{P}$-decomposition of $x\in\mathcal{R}^{n}$ as the $(m+1)$-tuple $\pi(x)=(\pi_{0}(x),\ldots,\pi_{m}(x))$ where, for each $i\in\{0,\ldots,m\}$,
(31) $\pi_{i}(x)=\left|\,{\left\{{k\in\{1,\ldots,n\}\;|\;x_{k}\in I_{i}}\right\}}\,\right|.$
Additionally, we denote by $\mathbb{D}^{\operatorname{wt}}_{n}$ the set of all $\mathcal{P}$-decompositions. Hence, given the set $\mathbb{D}^{\operatorname{wt}}_{n}=\left\{{\pi^{(0)},\ldots\pi^{(D-1)}}\right\}$ for some positive integer $D$, the induced symmetrized partition $\mathcal{P}_{\mathrm{sym},n}$ is given by
(32) $\mathcal{P}_{\mathrm{sym},n}=P^{\operatorname{wt}}_{\pi^{(0)}}\;|\;\cdots\;|\;P^{\operatorname{wt}}_{\pi^{(D-1)}}.$
That is, for every $i\in\left\{{0,\ldots,D-1}\right\}$,
(33) $P_{\pi^{(i)}}^{\operatorname{wt}}\coloneqq\left\{{x\in\mathcal{R}^{n}\;|\;\pi(x)=\pi^{(i)}}\right\}$
counts the number of $x\in\mathcal{R}^{n}$ having $\mathcal{P}$-decomposition $\pi^{(i)}$. Different to the original MacWilliams identities for the Hamming metric, we aim at enumerating all codewords of the same $\mathcal{P}$-decomposition $\pi^{(i)}\in\mathbb{D}^{\operatorname{wt}}_{n}$. That is, for each $\pi^{(i)}\in\mathbb{D}^{\operatorname{wt}}_{n}$, we are interested in
(34) $\mathcal{D}^{\operatorname{wt}}_{\pi^{(i)}}(\mathcal{C})\coloneqq\left|\,{P^{\operatorname{wt}}_{\pi^{(i)}}\cap\mathcal{C}}\,\right|.$
Furthermore, we denote the symmetrized partition enumerator by
(35) $\mathcal{D}^{\operatorname{wt}}(\mathcal{C}):=\left(\mathcal{D}^{\operatorname{wt}}_{\pi^{(0)}}(\mathcal{C}),\ldots,\mathcal{D}^{\operatorname{wt}}_{\pi^{(D-1)}}(\mathcal{C})\right).$
If the code $\mathcal{C}$ considered is clear from the context, we will simply write $\mathcal{D}^{\operatorname{wt}}$ and $\mathcal{D}^{\operatorname{wt}}_{\pi^{(i)}}$.
Given two $\mathcal{P}$-decompositions $\pi,\rho\in\mathcal{D}^{\operatorname{wt}}$, with the aim of computing the component-wise product of two tuples with decompositions $\pi=(\pi_{0},\ldots,\pi_{m})$ and $\rho=(\rho_{0},\ldots,\rho_{m})$, respectively, we denote the set of all compositions of $\rho$ with respect to $\pi$ by
(36) $\mathrm{Comp}^{\operatorname{wt}}_{\pi}(\rho):=\Big\{t=(t_{0},\ldots,t_{m})\,\Big|\,$ $t_{i}=(t_{i0},\ldots,t_{im})\in\left\{{0,\ldots,\pi_{i}}\right\}^{m}\text{ with }$
(37) $\sum_{j=0}^{m}t_{ij}=\pi_{i}\text{ and }\sum_{i=0}^{m}t_{ij}=\rho_{j},\text{for all }i,j\in\left\{{0,\ldots,m}\right\}\Big\}.$
Figure 1 illustrates the set $\mathrm{Comp}^{\operatorname{wt}}_{\pi}(\rho)$, i.e., the compositions of $\rho\in\mathbb{D}^{\operatorname{wt}}_{n}$ with respect to $\pi\in\mathbb{D}^{\operatorname{wt}}_{n}$.
[…]
Figure 1. Illustration of $\mathrm{Comp}^{\operatorname{wt}}_{\pi}(\rho)$.
In the following subsections we focus on the Lee, homogeneous and subfield metric. For each of them, we define an induced symmetrized partition using the above notation and show that this partition is Fourier-reflexive and allows to recover the weight enumerator.
3.1. Symmetrized Lee Partition for Additive Weights
Although the MacWilliams identities for the symmetrized Lee partition enumerator are well-established [23, 18, 22], in this section we propose an alternative, yet equivalent, formulation of the MacWilliams-type identities for the Lee decomposition enumerator presented in [18, Theorem 3.5]. Specifically, we explicitly calculate the Krawtchouk coefficients, providing a set of linear equations that enable us to recover the Lee decomposition enumerator as a function of the Lee decomposition enumerator of its dual, for any linear code over any finite chain ring and any additive weight.
In this section, if the characteristic of the ring is different from 2, we label the $q^{s}$ elements of $\mathcal{R}$ as follows:
(38) $\mathcal{R}=\{\alpha_{0}=0,\alpha_{1},\dots,\alpha_{M},\alpha_{M+1},\dots,\alpha_{q^{s}-1}\},$
@@ 3.1. Symmetrized Lee Partition for Additive Weights
where $M\coloneqq\left\lfloor{q^{s}/2}\right\rfloor$ and $\alpha_{q^{s}-i}$ represents the additive inverse of $\alpha_{i}$, i.e., $\alpha_{q^{s}-i}=-\alpha_{i}$. We refer to $\alpha_{q^{s}-i}$ as the opposite of $\alpha_{i}$. To enable the recovery of the Lee weight enumerator, we introduce the Lee partition of $\mathcal{R}$, in which each block consists of an element and its opposite.
@@
Given a finite chain ring $\mathcal{R}$ of characteristic greater than 2, for any $0\leq i\leq M$ we set $I^{\mathsf{L}}_{i}\coloneqq\{\alpha_{i},-\alpha_{i}\}$ to be the subset of $\mathcal{R}$ containing $\alpha_{i}$ and its opposite. We then define the Lee partition of $\mathcal{R}$ as
(39) $\mathcal{P}^{\mathsf{L}}=I^{\mathsf{L}}_{0}\;|\;I^{\mathsf{L}}_{1}\;|\;\cdots\;|\;I^{\mathsf{L}}_{M}\ .$
If the characteristic of $\mathcal{R}$ is 2, then $\left|\,{\mathcal{R}}\,\right|=2^{r}$ and every element is its own additive inverse, meaning $\alpha_{i}=-\alpha_{i}$ for all $i$. In this case, there is no need for relabeling, and the Lee partition
(40) $\mathcal{P}^{\mathsf{L}}=I^{\mathsf{L}}_{0}\;|\;I^{\mathsf{L}}_{1}\;|\;\cdots\;|\;I^{\mathsf{L}}_{2^{r}}\ .$
consists of the singletons, that is $I_{i}^{\mathsf{L}}=\left\{{\alpha_{i}}\right\}$.
Following the notation introduced at the beginning of the section, we denote by $\mathcal{P}_{\mathrm{sym},n}^{\mathsf{L}}$ the induced symmetrized Lee partition of $\mathcal{R}^{n}$.
Proposition 23.
The symmetrized Lee partition $\mathcal{P}_{\mathrm{sym},n}^{\mathsf{L}}$ is Fourier-reflexive.
Proof.
According to [18, Theorem 3.3], if contains $\{0\}$ as a block and $\mathcal{P}^{\mathsf{L}}$ is Fourier-reflexive, then $\mathcal{P}_{\mathrm{sym},n}^{\mathsf{L}}$ is Fourier-reflexive. Therefore, it suffices to prove that $\mathcal{P}^{\mathsf{L}}$ is Fourier-reflexive. If the characteristic of the ring is greater than 2, the Lee partition can be interpreted as the partition induced by the orbits of the action of the group $\{\pm 1\}$ on $\mathcal{R}$. It is shown in [18, Theorem 2.6] that any partition of a finite group induced by the orbits of a subgroup of its automorphism group is Fourier-reflexive. On the other hand, if the characteristic of the ring is 2, then the Lee partition consists of singleton blocks, which is clearly a Fourier-reflexive partition [5, 18]. Hence, the claim follows. ∎
Corollary 24 (MacWilliams-type Identities for the Symmetrized Lee Weight Enumerator).
Let $\mathcal{R}$ be a finite chain ring, with residue field size $q=p^{r}$ and nilpotency index $s$. For a positive integer $n$, consider a linear code $\mathcal{C}\subseteq\mathcal{R}^{n}$ and let $\mathcal{C}^{\perp}$ denote its dual. For a given Lee decomposition $\rho\in\mathbb{D}^{\mathsf{L}}_{n}$, the number of codewords in the dual $\mathcal{C}^{\perp}$ of Lee decomposition $\rho$ is given by
(41) $\mathcal{D}_{\rho}^{\mathsf{L}}(\mathcal{C}^{\perp})=\frac{1}{\left|\,{\mathcal{C}}\,\right|}\sum_{\pi\in\mathbb{D}^{\mathsf{L}}_{n}}K^{\mathsf{L}}_{\pi}(\rho)\mathcal{D}_{\pi}^{\mathsf{L}}(\mathcal{C}).$
Here, $K^{\mathsf{L}}_{\pi}(\rho)$ denotes the value of the Krawtchouk coefficient of $\rho$ with respect to $\pi$ and, for a given $p^{s}$-th root of unity $\xi$, it is given by
(42) $K^{\mathsf{L}}_{\pi}(\rho)\coloneqq\sum_{t\in\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)}\Bigg(\prod_{i=0}^{M}\binom{\pi_{i}}{t_{i0},\ldots,t_{iM}}\kappa^{\mathsf{L}}\Bigg),$
where